OpenStudy (anonymous):

Find the area of a rectangle with the following measurements: length = 2a2 + 3a -6 width = a + 2 A. 2a^3 + 7a^2 - 12 B. 2a^3 + 7a^2 - 12a - 12 C. 2a^3 + 3a^2 - 6a D.3a^3 + 7a^2 -12

OpenStudy (unklerhaukus):

area of a rectangle is length times width \(A = \ell \times w \) \( = ( 2a^2 + 3a -6) \times (a + 2) \) \( = \)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

A=L*W =(2A square + 3A -6) * (a+2) =(2A square +3A square -6A )(4A square +6A -12) ( Arrange there term with cross - and +) =Here you go on solving

OpenStudy (anonymous):

and then you will get answer and match with your question

OpenStudy (anonymous):

ok thank you again lol

OpenStudy (anonymous):

you well come B

OpenStudy (anonymous):

better lol

OpenStudy (anonymous):

haha lol

OpenStudy (anonymous):

cya soon

OpenStudy (anonymous):

oh yeah see